OpenStudy (anonymous):

MEDAL What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

The equation for the slope of a line is: (y2-y1)/(x2-x1) in this case: (8-(2))/((-5)-(-3)) = (6)/(-2) = -3 so we know our line will have the form y = 3x + c (c is a constant equal to the y intercept) to find c we plug in one of your 2 points, lets take (-3,2) therefore: 2 = (-3)(-3) + c 2 = 9 + c c = -7 hence, y = (-3)x - 7 or 3x+y = -7 Have a nice day!